OpenStudy (anonymous):

Find the lateral area the regular pyramid. with a slant of 4in. and base sides are 4in

OpenStudy (anonymous):



OpenStudy (anonymous):

Pleas help

OpenStudy (anonymous):

Is base a triangle?

OpenStudy (anonymous):

no sorry lol

OpenStudy (anonymous):

There are 4 triangle facets. I think \[Area = 4 \times 4 \times 4 \times \sin60\] \[Area = 32\sqrt{2}\]

OpenStudy (anonymous):

Each triangle has an area \[8\sqrt{2}\]

OpenStudy (anonymous):

was that the L.A. or T.A.?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

lateral area will be 4 times the area of triangle.

OpenStudy (anonymous):

32 square root 8?

OpenStudy (anonymous):

Silon?

OpenStudy (anonymous):

oops sorry.\[\sin60 = \frac{ \sqrt{3} }{ 2 }\] The area is \[32\sqrt{3}\]

OpenStudy (anonymous):

OH ok so the L.A. is \[\sqrt[32]{3} \]

OpenStudy (anonymous):

hehe, I mean \[32\times \sqrt{3}\]

OpenStudy (anonymous):

ok? i am looking for an answer that looks like _ _ \[\sqrt{_}\]

OpenStudy (anonymous):

sorry im new to this the answer should look like _ _ sqr _

OpenStudy (anonymous):

32 sqrt 3 ?

OpenStudy (anonymous):

no that answer is only partially right

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

partially right?

OpenStudy (anonymous):

yes i am doing the question online and it say it is only partially correct

OpenStudy (anonymous):

Can you send me print screen pic as a attachment, so I can see?

OpenStudy (anonymous):

https://media.glynlyon.com/g_geo_2012/8/groupi77.gif

OpenStudy (anonymous):

did that help?

OpenStudy (anonymous):

Got it, Actually, we have 3 facets. The base is a triangle. \[LA = 3 \times 4 \times 4 \sin60 = 24\sqrt{3}\]

OpenStudy (anonymous):

awesome thank you SO MUCH. can you also find the T.A.?

OpenStudy (anonymous):

What is TA?

OpenStudy (anonymous):

triangle area?

OpenStudy (anonymous):

Total Area

OpenStudy (anonymous):

Sure, \[TA = 4 \times 4 \times 4 \sin60 = 32\sqrt{3}\]

OpenStudy (anonymous):

You awesome. now if your up for it? can you find the volume?

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

oh no and i just submitted those answers and they are still partially correct.

OpenStudy (anonymous):

Let me correct mistakes.each triangle has an area \[Area of \triangle = \frac{ 4\times4 \times \sin60 }{ 2 }\]

OpenStudy (anonymous):

Area of triangle is 4 sqrt 3

OpenStudy (anonymous):

\[LA = 3\times4\sqrt{3} =12\sqrt{3}\]

OpenStudy (anonymous):

\[TA = 4\times 4\sqrt{3} = 16\sqrt{3}\]

OpenStudy (anonymous):

\[The Volume =\frac{ 4^{3} }{ 6\sqrt{2} } = \frac{ 16\sqrt{2} }{ 3 }\]

OpenStudy (anonymous):

Sorry that I made a stupid mistake :(

OpenStudy (anonymous):

Hello?

OpenStudy (anonymous):

hey sorry i thought you left.

OpenStudy (anonymous):

All of those answers are correct thank you so much.

OpenStudy (anonymous):

You are welcome :)